Fermi National Laboratory
 

Accelerator Update for April 3, 2009 - April 6, 2009

 

The Integrated Luminosity 3/30/09 to 4/6/09 was approximately 56.72 Pb-1.

Friday April 3

The day shift began with Operations monitoring the Tevatron (TeV) store 6955 (203.73E30) and the Recycler stash (110.7E10), with the Antiproton Source stacking (13.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

The NuMI Spokesperson reported at 8:31 AM that the experiment was beginning a special run mode with the NuMI horn off. This run will last two to three days.

The Neutron Therapy Facility (NTF) reported at 11:37 AM that they had finished treating their last patient for the day.

At 1:24 PM, the P1 transfer line beam permit tripped off when a power supply’s (V714) current drifted low. Everything reset.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons eight times during the day shift.

At 5:02 PM, a Linac RF station (LRF3) tripped off. It reset.

Operations terminated store 6955 at 8:15 PM.

Operations put the TeV into shot setup at 8:26 PM.

Operations established store 6956 at 9:50 PM with an initial luminosity of 320.64E30.

Operations reported at 11:57 PM that the Recycler had stashed seven times during the evening shift.

Saturday April 4

The midnight shift began with Operations monitoring store 6956 (213.23E30) and the stash (113.4E10), with Pbar stacking (15.1E10), and with NuMI and MiniBooNE taking beam.

At 4:29 AM, LRF2 tripped off. It reset.

At 7:26 AM, LRF3 tripped off. It reset.

Operations reported at 8:02 AM that the Recycler had stashed eight times during the midnight shift.

Operations terminated store 6956 at 11:54 AM.

Operations put the TeV into shot setup at 12:09 PM.

At 1:17 PM, cryo technicians at CHL reported that they were having cold box turbine problems. Operations put shot setup on hold.

At 2:03 PM, the Cryo Coordinator reported that they would have to warm up cold box #2, which means they would have to warm up the TeV to nitrogen temperature. They have to replace the cold box #2 turbine. Operations aborted the protons and antiprotons in the TeV.

Operations reported at 2:59 PM that they had closed all TeV beam valves that could be closed remotely.

At 5:16 PM, Operations reported that the Pbar Lithium lens water return conductivity had caused an interlock trip. It reset.

Operations began preparing the TeV F sector for access at 5:33 PM.

At 5:45 PM, the lithium lens tripped off again. Operations contacted a Pbar Physicist.

At 6:26 PM, a Pbar Physicist reported a lens failure, probably due a LCW leak. Operators transferred the antiprotons to the Recycler and put Pbar into standby. The lens will be replaced on Monday.

Operations reported at 8:18 PM that the Recycler had stashed nine times during the last 12 hours.

At 8:21 PM, Operations received a high temperature alarm for two MI-30 magnets. Operations halted the ramp and investigated.

The Linac Group Leader reported at 8:39 PM that she was running study pulses.

An MI Operations Specialist reported at 8:51 PM that operators should check the PLC and call in EE Support. Operators did this.

The Proton Source Operations Specialist reported at 8:54 PM that he had put the Booster into dump mode and was looking at the Linac beam coming into Booster.

The EE Support person reported at 10:49 PM that he had replaced a 24-volt power supply for the PLC. MI resumed ramping.

The TeV F sector access ended at 10:59 PM.

Operations established beam to NuMI and MiniBooNE at 11:53 PM.

Sunday April 5

The midnight shift began with cryo system technicians warming the TeV to replace a cold box turbine, with Operations monitoring the stash, with Pbar off until Monday due to a lithium lens failure, and with NuMI and MiniBooNE taking beam.

At 12:05 AM, MIRF6 tripped off for the first of two times in 2.5 hours. It reset.

At 1:02 AM, LRF3 tripped off. It reset.

At 5:54 AM, the Booster East Anode Power Supply tripped off. It reset.

Operations prepared the TeV F sector for an access at 11:17 AM.

At 11:44 AM, Operations reported that LRF3 had tripped off four times in 10 minutes. Operators investigated and the station decided to operate stably.

Recycler Physicists began a study period at 12:29 PM.

The Cryo Coordinator reported at 3:45 PM that they had brought the backup cold box online (cold box #1). He expected that it would be ready to liquefy helium during the midnight shift.

The F sector access ended at 4:35 PM.

Operations established beam to NuMI and MiniBooNE by 5:02 PM.

At 5:36 PM, BRF15 and 16 tripped off due to low vacuum pressure. They both reset.

The Recycler study ended at 7:11 PM.

At 7:35 PM, Operations received a TeV sector C0 cold compressor room smoke detector alarm. Operators and the Fermi Fire Department investigated and discovered a compressor discharge oil line had broken and sprayed oil onto warm parts of the compressor, which caused the smoke. Operations turned off the compressor, valved off the oil, and called in an ES&H Waste Coordinator to manage the oil spray.

At 8:47 PM, Operations lowered the MI beam intensity due to beam position problems.

At 10:44 PM, Operations reported problems with the I- Source’s regulation.

At 11:59 PM, MIRF6 tripped off. It reset.

Monday April 6

The midnight shift began with the CHL cold box #1 cooling down and cold box #2 warming up, with Operations monitoring the stash, with Pbar in standby, and with NuMI and MiniBooNE taking beam.

CHL reported 1:03 AM that it was beginning to pump liquid helium to the TeV.

Operations began preparing Booster, TeV F sector, and MI and MI-12B for access at 6:01 AM.

Operations began preparing Pbar for access at 6:24 AM.

Booster RF technicians turned off the West RF system at 7:48 AM.

The Plans for Monday

The plans for today are to replace the lithium lens. The Cryo Coordinator reported that the TeV would be ready around 4 PM on Tuesday, 4/7/09. The Run Coordinator will allow Linac and Booster to conduct maintenance until noon today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/06/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory